find the magnification of an object placed 10cm from a convex miror of radius of curvature 20 cm .

find the  magnification of an object placed 10cm from a convex miror of radius of curvature 20 cm .

Grade:11

1 Answers

Arun
6 years ago
0.5 is the magnification

it came because......
u= -10cm
f=10 cm
v=5cm {1/f=1/v+1/u}

magnification=-v/u
m=5/10
magnification=0.5

hope it is correct